OpenStudy (anonymous):

Algebra help! Inequality!!! 1. The product of a number and 8 is at least 25. 2.The quotient of a number and 12 is no more than 6.

OpenStudy (anonymous):

number one 8*x>=25 x>=25/8=3.125 other one i dont know

OpenStudy (anonymous):

Thanks!!! :)

OpenStudy (anonymous):

yea other one does not make sense to me but i think it could be n/12 <= 6 but not sure at all

OpenStudy (anonymous):

Thanks so much @samstopp !!! :)

Directrix (directrix):

The quotient of a number and 12 is no more than 6. ------- If x represents the number, then the associated inequality is: x/12 ≤ to 6. No more than 6 means less than or equal to 6.

Directrix (directrix):

If an answer is required, solve for x to get x ≤ 72.